1. You conduct a one-factor ANOVA with 3 groups and 10 subjects in each group (a balanced design) and obtain F=3.28F=3.28. Use Excel to find the requested values (report accurate to 4 decimal places).
dfbetween=
dfwithin=
p=

2.You conduct a one-factor ANOVA with 3 groups and 7 subjects in each group (a balanced design). Use Excel to find the critical values for α=0.02α=0.02 and α=0.01α=0.01 (report accurate to 3 decimal places).
F0.02=
F0.01=
Answer #1

Ans:

1)N=3*10=30

df(between)=3-1=2

df(within)=30-3=27

p=fdist(3.28,2,27)=0.0531

2)N=3*7=21

df(between)=3-1=2

df(within)=21-3=18

F0.02=FINV(0.02,2,18)=4.900

F0.01=FINV(0.01,2,18)=6.013